OpenStudy (anonymous):

If x*x + 2x - 35 = 0, then x = ?

OpenStudy (anonymous):

is \(xx\) supposed to be \(x^2\) ?

OpenStudy (anonymous):

\[x^2+2x-35=0 \rightarrow (x+7)(x-5)=0 \implies x = -7,5\] @RAWFECTION

OpenStudy (mertsj):

@genius12 Please refrain from posting answers. Please try to lead the asker to the answer

OpenStudy (anonymous):

ok so u r not supposed to giv anybody the answer?

OpenStudy (anonymous):

@Mertsj I am sorry. Usually I try to tell and explain and hint to what is the answer but sometimes I forget the no giving answer rule. I'll watch out next time.

OpenStudy (mertsj):

We try to lead the student to the answer if at all possible.

OpenStudy (unklerhaukus):

you could use the quadratic formula \[x^2+ 2x - 35 = 0\qquad\implies\qquad x_{1,2}=\frac{-2\pm\sqrt{2^2-4\times1\times-35}}{2\times1} \]